1. The physician has recommended that a patient with pain from osteoarthritis take off excess pounds. The Dr. explains that being overweight increases the load put on the joints. The patient weighs 110 kg. This is ______lb.

2. Ordered: Ibuprophen 0.5 g
Have avaliable : Ibuprophen 250 mg tablets
Give: ____________tablets?

3. For dinner, the patient took 1 cup of milk, 3 Tbs of ice cream (this is counted as liquid intake), and a 12 oz can of soda. What is the patient intake for dinner? _____mL

4. A patient is prescribed 20 mg of ketorolac (Toradol) intramuscularly. How many milliliters should you draw up from a vial of ketorolac containing 30 mg/mL? (Note: Fluids should be rounded to the nearest 10th)

a. 0.5mL

b. 0.7mL

c. 1mL

d. 2mL

5. Which is the only NSAID recommended for children? _________

a. naproxen

b. celecoxib

c. aspirin

d. ibupophen
Answer #1

1.The patient’s weight= 110 kg. = 110x 2.2 lb= 242 lb

2. Ordered ibuprophen= 0.5 g= 500mg

Available ibuprophen =250 mg tablets

Give: ____________tablets?

500/250= 2 tablets.

For dinner, the patient took 1 cup of milk= 240 ml

3 Tbs of ice cream (this is counted as liquid intake)= 45 ml

As 1 tbs= 15 ml

12 oz can of soda= 360 ml

As 1 oz = 30 ml

What is the patient intake for dinner? _____mL

Total = 240+45+360= 645 ml

4.A patient is prescribed 20 mg of ketorolac (Toradol) intramuscularly.

Available ketorolac =30 mg/mL

The required ml to be withdrawn to fulfill the order of 20 mg= 20/30= 0.666

= 0.7 ml

Hence option b is correct.

5. Option D is correct as

Ibuprofen is the NSAID which is considered safe in children of 6 months and older